A ray of light in air “n1”is incident strikes with an angle of incidence of 60 degrees on flint glass”n2”. Calculate the angle of refraction to the nearest degree (8 pts).


1
Expert's answer
2021-07-01T04:58:01-0400
"\sin{\theta}=\frac{\sin{60}}{1.69}\\
\theta=\arcsin{\frac{\sin{60}}{1.69}}=30.8\degree"